-2(z+3)-z=-z-4(z+2)
Answer by stanbon(75887)

-2(z+3)-z=-z-4(z+2)
-------------
-2z-6-z = -z-4z-8
-----
-3z-6 = -5z-8
-----
2z = -2
z = -1
------------------
Cheers,
Stan H.
------------